HETAV PATEL

Grade 11,

The resistance of a wire is 10 ohm. It`s length is increased by 10% by stretching. The new resistance will be(1)12ohm (2) 1.2ohm (3)1.3ohm (4)1.1ohm

The resistance of a wire is 10 ohm. It`s length is increased by 10% by stretching. The new resistance will be(1)12ohm (2) 1.2ohm (3)1.3ohm (4)1.1ohm

Grade:11

1 Answers

Vikas TU
7 years ago
As Resistance is directly proportional to the length of the resistance wire,
R = kl^2
After length gets increased by 10% we get,
R’ = k(l + 0.l)^2
divide both we get,
R/R’ = 1/(1.1)^2 => 1/1.21
R’ = 1.21 R ohm
that is second option.